OpenStudy (anonymous):

the total cost of an item includes sales tax is directly proportional to its price. if the total cost of a $24 item is $25.44 , what is the total cost on a $75 item

OpenStudy (anonymous):

25.44/24 = x/75 75*25.44 = 24x x = 75*25.44/24 = 79.5

OpenStudy (anonymous):

What did you get?

OpenStudy (triciaal):

directly proportional means as one goes up/down the other goes up/down by a certain constant factor.